### v1

@@ 3 Truthfulness under Completely Overlapping Ownership
Now returning to our more general setup with the overlapping ownership, to aggregate information from multiple owners, the most natural design is perhaps to use the (weighted) averaged scores from the quality estimates based on each owner’s reported ranking. This leads to our extension of the Isotonic Mechanism [23] for the multi-owners setup, as presented in Mechanism 1. In this mechanism, one can simply set the weight parameter $\alpha^{j}=1/m$ to evenly weigh on each owner’s reported ranking, or set any $\alpha^{j}\in[0,1]$ for each owner $j$ (e.g., due to their expert levels, track records) as a convex combination of scores calibrated by each owner’s reported ranking — we leave such kind of implementation choices to the practitioner’s discretion.
Input : Review scores $\bm{y}\in\mathbb{R}^{n}$, Weights $\{\alpha^{j}\}_{j=1}^{m}.$
for  $j\in[m]=\{1,2,\dots,m\}$  do
@@
Set $\widehat{\bm{R}}\leftarrow\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}^{j}$.
@@
In this section, we consider the setting where every submission has the same set of authors, referred to as the special case of completely overlapping ownership. As it turns out, truth-telling forms a Nash equilibrium in this setting under Mechanism 1444This result can be related to the ex-post incentive compatibility of Mechanism 1, a common goal of truthful mechanism design. Moreover, this equilibrium is payoff dominant [11], i.e., one that is Pareto superior to all other Nash equilibria in the game. Put simply, all owners would prefer this equilibrium, because it simultaneously gives every owner the highest equilibrium utility among all possible equilibrium outcomes. This nice property makes it more plausible to expect agents’ truthful behaviors, despite the potential existence of multiple equilibria in the game; more discussions about this mechanism’s truthfulness from the perspective of behavioral theory can be found in Remark 1.1.
@@
Remark 1.1 (Additional Discussions on the Strategic Game Induced by the Isotonic Mechanism).
The strategic game induced by Mechanism 1 features a certain structure that is known as the common interest game in the literature [11, 4]. We briefly highlight this connection here and discuss evidence from both empirical human preferences and behavioral theory about why truthful behavior should be expected in such games. In its simplest form, the payoff matrix of a common interest game typically takes one of the three structures illustrated in Table 1 (only orders of the payoff values matter). Namely, there is an action profile (the upper left cell of each game) that represents the common interest of both players; this corresponds to the truthful action profile $(\pi^{\star},\pi^{\star})$ under the Isotonic Mechanism which forms a payoff-dominant Nash equilibrium, according to Theorem 1. Meanwhile, the three variants of common interest games capture the different possible orders of payoff values in the lower right cell and those in the off-diagonal cells. Any of these payoff variants may be realized under the Isotonic Mechanism, since our proof using the Jensen’s inequality only offers upper bounds on the utilities of these non-truthful action profiles (see e.g., Equation (4)).
@@
3.1 Non-truthfulness Beyond Completely Overlapping Ownership.
Despite the stability of truthful NE shown in Theorem 1, the ownership relation is not complete in many application scenarios, e.g., different papers are written by different sets of authors. A natural extension of Mechanism 1 for the case of incompletely overlapping ownership is to average the scores according to the ownership, $\widehat{R}_{i}\leftarrow{\sum_{j=1}^{m}e^{j}_{i}\widehat{R}^{j}_{i}}/{\sum_{j^{\prime}=1}^{m}e^{j^{\prime}}_{i}}$, where the binary ownership indicator $e^{j}_{i}=1$ if and only if author $j$ owns item $i$. The question is whether truthful reporting still forms a Nash equilibrium. Unfortunately, the answer turns out to be “No”. Specifically, the owners may gain from the misreporting strategy and one possible manipulation is to use the reputation of good items to promote a bad item — it may only hurt the good item a little, but help the bad item a lot. We demonstrate such kind of strategic manipulations via a concrete example below.
Example 1 (Non-truthfulness of the Isotonic Mechanism under Partially Overlapping Ownership).
Consider a case of $m$ owners and $n=3$ items. The items have ground-truth qualities $R_{1}=9,R_{2}=8,R_{3}=4$ — one weak and two strong. The ownership is not completely overlapping. The first $m-1$ owners work together on the two strong items $1,2$, while the $m$-th owner works on the item $2,3$. Let the $j$-th owners’ utility be $\mathrm{U}^{j}(\widehat{\bm{R}})=\sum_{i\in\mathcal{I}^{j}}\mathop{max}\{R_{i}-5,0\}$. For simplicity, suppose the reviews are noiseless so that, if all owners report their ranking truthfully, the calibration is perfect, i.e., $\widehat{\bm{R}}=\textbf{y}=\bm{R}$. See Table 2 for an illustration.
@@
Item Owner $1$ $2$ $m-1$ $m$
1 1 1 1 0
2 1 1 1 1
3 0 0 0 1
Table 2: An illustration of item scores and the ownership matrix $E=(e_{i}^{j})_{m\times n}$ in Example 1
However, the $m$-th owner does not have the incentive to report the ranking of items $1,2$ truthfully, given that the first $m-1$ owners report the true ranking. That is, under the flipped ranking that $\widetilde{R}_{1}^{m}\geq\widetilde{R}_{2}^{m}$, Mechanism 1 would have $\widetilde{R}_{1}^{m}=\widetilde{R}_{2}^{m}=6$ and averaged calibrated scores $\widetilde{R}_{1}=9,\widetilde{R}_{2}=\widetilde{R}_{2}^{m}+\sum_{j=1}^{m-1}R_{2}^{j}=8-\frac{2}{m},\widetilde{R}_{3}=\widetilde{R}_{3}^{m}=6$. We can see that the utility of $m$-th owner under truthful ranking is, $13<3-\frac{2}{m}+$ 1, strictly worse than that under the non-truthful ranking for any $m\geq 3$.
At a high level, such strategic behavior is due to the owner’s uneven influence on the average scores of its different items. One may speculate that we can potentially resolve this kind of situation through a careful reweighing process of each owner’s influence. The answer also turns out to be “No”. Specifically, in the example above, $m$-th owner, as the sole owner of item 1, would always fully determine the score of item 1; yet, this owner can never fully determine the score of item 2 unless we choose to ignore other owners’ opinion or assign some weight to the raw review scores, which more or less defeats the purpose of this community effort of calibrating review scores. Hence, we need to seek a different approach to help us truthfully aggregate the authors’ information in practice.
@@ 4 Restoring Truthfulness via Partitioning
To ensure the truthfulness of the Isotonic Mechanism beyond complete ownership, this section studies a partition-based approach. Our main idea is to partition the ownership sets into multiple blocks, ensuring complete overlap within each block by some owners. We then apply the Isotonic Mechanism separately to each block, eliciting truthful rankings only from the owners who completely own that block. For example, this includes authors who wrote all the papers within that block (refer to the illustrations in Figure 2 and Table 3). Finally, we use the elicited rankings from these blocks to estimate the item quality. This generalized approach is described in the following Mechanism 2.
Input : Review score $\bm{y}\in\mathbb{R}^{n}$.
Input : Any partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ of all items (i.e., $\bigcup_{k=1}^{K}{\mathcal{S}}_{k}=[n]$ and ${\mathcal{S}}_{k}\cap{\mathcal{S}}_{k^{\prime}}=\emptyset$).
for every ${\mathcal{S}}_{k}\in{\mathcal{S}}$ do
@@
             Apply Mechanism 1 to item set ${\mathcal{S}}_{k}$ with owner set ${\mathcal{T}}_{k}$ to estimate the quality of items in ${\mathcal{S}}_{k}$, denoted by $\widehat{\bm{R}}[{\mathcal{S}}_{k}]$.
@@
We remark that in Mechanism 2, each owner $j\in{\mathcal{T}}_{k}$ has the complete ownership of the items in ${\mathcal{S}}_{k}$ by construction. As a corollary of Theorem 1, it is easy to show that Mechanism 2 is truthful under any partition ${\mathcal{S}}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ (which is part of the input to Mechanism 2), because we only elicit rankings of blocks from those who completely own the block. Thus, all owners $j\in{\mathcal{T}}_{k}$ will report truthful ranking over items in ${\mathcal{S}}_{k}\subseteq\mathcal{I}^{j}$. Owing to the independence of quality estimates for items between different partition blocks, the overall mechanism is truthful; we formalize it in the following corollary.
@@
Under Mechanism 2 with any item set partition, it forms a Bayes-Nash equilibrium for each owner to truthfully report the ranking of items within each block specified by the partition.
Remark 1.2 (Practical Implementation of Mechanism 2).
@@
While Mechanism 2 only elicits ranking of items within each block ${\mathcal{S}}_{k}$, in practical implementations, the mechanism designer may simply ask the owners to rank all their items but commit to only use part of each owner’s ranking as specified by the mechanism. Under such a committed mechanism, any owner will be indifferent about the order of any two items across any two different partition sets ${\mathcal{S}}_{k}$ and ${\mathcal{S}}_{k^{\prime}}$ since their order will never be used by the mechanism. Thus, the owner can be assumed to break ties in favor of the designer and reveal their full ranking truthfully. [^5] There is also a more game-theoretic way to justify such tie-breaking by adding a negligible amount of randomness for picking an arbitrarily different partition, which consequently creates a negligible amount of incentive for any owner to report other rankings truthfully. Such equilibrium refinement via randomization has appeared in mechanism design literature [18]. Notably, however, the designer’s commitment to not using any order information beyond what the mechanism specified (despite such information being elicited) is important since otherwise, truthfulness will not hold. In reality, the mechanism designer as a trusted authority (e.g., the organizers of a large ML conference) has such commitment power.
[Uncaptioned image] Figure 2: A partition of partially overlapping ownership Item Owner $1$ $2$ $3$ $4$ $1$ 1 1 1 0 $2$ 1 1 1 0 $3$ 1 1 1 0 $4$ 0 0 1 0 $5$ 0 0 1 1 $6$ 0 0 1 1 Table 3: A visualization via table
While any partition will lead to a truthful Mechanism 2, what partition we choose will significantly affect how much information we can elicit from owners to refine our quality estimation. Intuitively, the smaller the size of a block is, the less ranking information we can elicit. [^6] For instance, the ranking of a size-$10$ item block has $\frac{10\times 9}{2}=45$ pairs of comparison, whereas the rankings of two size-$5$ item blocks have $202\times\frac{5\times 4}{2}=$ 20 pairs of comparison. Thus, the core of our partition-based approach is to design an efficient algorithm to compute high-quality partition structure, which we now elaborate on.

### v3

@@ 3 An Isotonic Mechanism for Completely Overlapping Ownership
In the general setup with overlapping ownership, we would like to design a mechanism that aggregates information from as many owners as possible, while preserving the desirable properties of truthfulness and statistical efficiency. The most natural design is perhaps to use the (weighted) averaged scores from the estimates based on each owner’s reported ranking. The mechanism takes the input of a problem instance, specified by review scores $\bm{y}\in\mathbb{R}^{n}$ and owner credentials $\{\alpha^{j}\}_{j=1}^{m}\in[0,1]^{m}$, and outputs the adjusted review score $\widehat{\bm{R}}$ using rankings elicited from all owners. Notably, the owner credential is a set of weights $\{\alpha^{j}\}_{j=1}^{m}$ that pre-specifies the different levels of influence of the owners; we view them as part of the problem instance that reflects each owner’s expert level, reputation scores and track records in the given instance. In practice, one could simply set $\alpha^{j}=1$ to evenly weigh on each owner’s reported ranking, or set personalized $\alpha^{j}$ for each owner $j$ to account for different owner’s expert level — we include it as a part of the input so that it is a choice to the practitioner’s discretion. We formally describe these procedures in Mechanism 1. The rest of this section is to analyze the situations when it does or does not work.
Input: Review scores $\bm{y}\in\mathbb{R}^{n}$, owner credentials $\{\alpha^{j}\}_{j=1}^{m}$.
for every $j\in[m]=\{1,2,\dots,m\}$  do
@@
Set $\widehat{\bm{R}}\leftarrow\sum_{j=1}^{m}\alpha^{j}\widehat{\bm{R}}^{j}/\sum_{j=1}^{m}\alpha^{j}$.
@@
3.1 Truthfulness under Completely Overlapping Ownership
We start by considering a useful special case in which every submission is owned by every owner, referred to the completely overlapping ownership situation. As it turns out, truth-telling forms a Nash equilibrium in this setting under Mechanism 1999This result can be related to the ex-post incentive compatibility of Mechanism 1, a common goal of truthful mechanism design. Moreover, this equilibrium is payoff dominant [16], i.e., one that is Pareto superior to all other Nash equilibria in the game. Put simply, all owners would prefer this equilibrium, because it simultaneously gives every owner the highest equilibrium utility among all possible equilibrium outcomes. This nice property makes it more plausible to expect agents’ truthful behaviors, despite the potential existence of multiple equilibria in the game; more discussions about the truthfulness of this mechanism from the perspective of behavioral theory can be found in Remark 1.
@@ 3.1 Truthfulness under Completely Overlapping Ownership
We defer the proof of Theorem 1 to Appendix A and conclude this subsection with additional evidence on the truthfulness of the Isotonic Mechanism according to the behavioral game theory.
Remark 1 (Truthfulness of Isotonic Mechanism from a behavioral angle).
Mechanism 1 induces the strategic game with special structures known as the common interest games [16, 2]. We briefly highlight the connections here and discuss evidence from both empirical human preferences and behavioral theory on why truthful behavior should be expected in such games. In its simplest form, the payoff matrix of a common interest game typically takes one of the three structures illustrated in Table 1 (only orders of the payoff values matter). Namely, there is an action profile (the upper left cell of each game) that represents the common interest of both players; this corresponds to the truthful action profile $(\pi^{\star},\pi^{\star})$ under the Isotonic Mechanism which forms a payoff-dominant Nash equilibrium, according to Theorem 1. Meanwhile, the three variants of common interest games capture the different possible orders of payoff values in the lower right cell and those in the off-diagonal cells. Any variant of these payoff matrices may be realized under the Isotonic Mechanism, since our proof using Jensen’s inequality only offers upper bounds on the utilities of these non-truthful action profiles (see e.g., Equation (5)).
@@
3.2 Non-truthfulness Beyond Completely Overlapping Ownership
Despite the nice properties of the Nash equilibrium shown in Theorem 1, its assumption of completely overlapping ownership is usually not satisfied in many application scenarios — different papers are often written by different sets of authors. To extend Mechanism 1 to the case of incompletely overlapping ownership, one natural choice is to average the scores according to the ownership, $\widehat{R}_{i}\leftarrow{\sum_{j=1}^{m}e^{j}_{i}\widehat{R}^{j}_{i}}/{\sum_{j^{\prime}=1}^{m}e^{j^{\prime}}_{i}}$, where the binary ownership indicator $e^{j}_{i}=1$ if and only if owner $j$ owns item $i$. The question is whether truthful reporting still forms a Nash equilibrium. Unfortunately, the answer turns out to be “No”. Specifically, the owners may gain from the misreporting strategy and one possible manipulation is to use the reputation of good items to promote a bad item — it may only hurt the good item a little, but help the bad item a lot. We demonstrate such kind of strategic manipulations via a concrete example below.
Example 1 (Non-truthfulness under Partially Overlapping Ownership).
Consider a case of $m$ owners and $n=3$ items. The items have ground-truth scores $R_{1}=9,R_{2}=8,R_{3}=4$ — one weak and two strong. The ownership is not completely overlapping. As illustrated in Table 2, the first $m-1$ owners work together on the two strong items $1,2$, while the $m$-th owner works on the item $2,3$. For each $j$, let the $j$-th owner’s utility be $\mathrm{U}^{j}(\widehat{\bm{R}})=\sum_{i\in\mathcal{I}^{j}}\mathop{max}\{R_{i}-5,0\}$.
Item Owner $1$ $2$cdots $m-1$ $m$
1 1 1 $\cdots$ 1 0
2 1 1 $\cdots$ 1 1
3 0 0 $\cdots$ 0 1
@@ 3.2 Non-truthfulness Beyond Completely Overlapping Ownership
Table 2: An illustration of the ownership matrix $E=(e_{i}^{j})_{m\times n}$ and ground-truth scores in Example 1.
For simplicity, suppose the reviews are noiseless so that, if all owners report their ranking truthfully, the calibration is perfect, i.e., $\widehat{\bm{R}}=\textbf{y}=\bm{R}$. However, we can observe that the $m$-th owner does not have the incentive to report the ranking of items $1,2$ truthfully, given that the first $m-1$ owners report the true ranking. That is, under the flipped ranking that $\widetilde{R}_{2}^{m}\leq\widetilde{R}_{3}^{m}$, Mechanism 1 would have $\widetilde{R}_{2}^{m}=\widetilde{R}_{3}^{m}=6$ and the adjusted scores $\widetilde{R}_{1}=9,\widetilde{R}_{2}=\frac{1}{m}(\widetilde{R}_{2}^{m}+\sum_{j=1}^{m-1}R_{2}^{j})=8-\frac{2}{m},\widetilde{R}_{3}=\widetilde{R}_{3}^{m}=6$. We can see that the utility of $m$-th owner under truthful ranking is, $13<3-\frac{2}{m}+$ 1, strictly worse than that under the non-truthful ranking for any $m\geq 3$.
At a high level, such strategic behavior is due to the owner’s uneven influence on the average scores of its different items — we will formalize this intuition in Theorem 2. One may speculate that we can potentially resolve this kind of situation through a careful reweighing process of each owner’s influence. The answer also turns out to be “No”. Specifically, in the example above, $m$-th owner, as the sole owner of item 3, would always fully determine the score of item 3; yet, this owner can never fully determine the score of item 2 unless we choose to ignore other owners’ opinion or assign some weight to the raw review scores, which more or less defeats the purpose of this community effort of calibrating review scores. Hence, we need to seek a different approach to help us truthfully aggregate the owners’ information in practice.
@@ 4 Restoring Truthfulness via Partitioning
To ensure the truthfulness beyond complete ownership, this section studies a partition-based approach. Our main idea is to partition the ownership sets into multiple blocks, ensuring complete overlap within each block by some owners. We then individually apply Mechanism 1 to each block, eliciting truthful rankings only from the owners who completely own that block, as illustrated in Figure 3. Finally, we use the elicited rankings from these blocks to estimate the ground-truth scores. We formally described this procedure in the following Mechanism 2.
Figure 3: A partition of partially overlapping ownership.
Parameters : Item set partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ based on ownership relations $\{\mathcal{I}^{j}\}_{j=1}^{m}$.
Input: Review score $\bm{y}\in\mathbb{R}^{n}$, owner credentials $\{\alpha^{j}\}_{j=1}^{m}$.
for every ${\mathcal{S}}_{k}\in\mathfrak{S}$ do
@@
             Apply Mechanism 1 to item set ${\mathcal{S}}_{k}$ with owner set ${\mathcal{T}}_{k}$ and weights $\{\alpha^{j}\}_{j\in{\mathcal{T}}_{k}}$ to estimate the ground-truth score of items in ${\mathcal{S}}_{k}$, denoted by $\widehat{\bm{R}}[{\mathcal{S}}_{k}]$.
@@
Compared to Mechanism 1, Mechanism 2 additionally determines a set of parameters which include a partition of item set $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ that satisfies $\bigcup_{k=1}^{K}{\mathcal{S}}_{k}=[n],{\mathcal{S}}_{k}\cap{\mathcal{S}}_{k^{\prime}}=\varnothing$. We ask the parameters to only depend on the ownership relations for practical considerations: since the mechanism may be truthful only under some proper choice of parameters, the owners should be informed of the parameters during submission, when the ownership relations are formed but the review scores are not realized. One can observe that in Mechanism 2, each owner $j\in{\mathcal{T}}_{k}$ has the complete ownership of the items in ${\mathcal{S}}_{k}$ by construction. As a corollary of Theorem 1, it is easy to show that Mechanism 2 is truthful under any partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ and any weights $\{\{\beta^{j}_{k}\}_{k=1}^{K}\}_{j=1}^{m}$, because we only elicit rankings of items within each block from those who completely own the block. Thus, all owners $j\in{\mathcal{T}}_{k}$ will report truthful ranking over items in ${\mathcal{S}}_{k}\subseteq\mathcal{I}^{j}$. Owing to the independence of score estimates for items between different partition blocks, the overall mechanism is truthful; we formalize it in the following corollary.
@@
Mechanism 2 is truthful for any input instance and parameter choice in the following sense: it forms a Bayes-Nash equilibrium for each owner to truthfully report the ranking of their items within each block specified by the partition.
Moreover, Mechanism 2 can be implemented in practice to promote both simplicity and fairness for participants. Below, we highlight several key aspects and advocate for field experiments to better evaluate and optimize for its practical applicability.
Remark 2 (Practical Implementation and Fairness Considerations).
@@
While Mechanism 2 only elicits the ranking of items within each block ${\mathcal{S}}_{k}$, the mechanism designer may implement the mechanism by simply asking owners to rank all their items with the commitment to only use part of the ranking of each owner as specified by the mechanism. Under such a committed mechanism, any owner will be indifferent about the order of any two items in any two different partition sets ${\mathcal{S}}_{k}$ and ${\mathcal{S}}_{k^{\prime}}$ since their order will never be used by the mechanism. Thus, the owner can be assumed to break ties in favor of the designer and reveal their full ranking truthfully. [^10] There is also a formal justification for such tie-breaking by adding a negligible amount of randomness for picking an arbitrarily different partition, which consequently creates a negligible amount of incentive for any owner to report other rankings truthfully. Such equilibrium refinement via randomization has appeared in the literature on mechanism design [29]. Notably, however, the designer’s commitment to not using any order information beyond what the mechanism specified (despite such information being elicited) is important, since otherwise, truthfulness will not hold. In reality, the mechanism designer as a trusted authority (e.g., the organizers of a large ML conference) usually has such commitment power.
When applying our mechanism to conference paper peer review, a potential limitation one might argue is that if an author only has one or two papers, then the mechanism may not elicit ranking information from this author, which may be “unfair” to this author’s papers. We remark that this is not a major concern because most papers in today’s ML conferences have multiple authors, so long as some author of a paper offered ranking information, that paper’s review score will be calibrated. For instance, in ICLR 2023, approximately 76.5% of papers had at least one “senior” author with multiple submissions and all these papers’ review scores could be calibrated based on these “senior” authors’ reported ranking. In addition, there are other ways to further mitigate potential unfairness. For example, while partitioning may lead to some owners’ information being unused across blocks, we could randomize the choice of partitions in the mechanism to ensure that each owner’s reported ranking is incorporated in the learning process with non-zero probability. That said, while our mechanism could mitigate the issue of fairness to some extent, we acknowledge that it does not come with a guarantee of perfect fairness which in general is an extremely deep concept, both socially and philosophically, and is also multifaceted. We expect follow-up research to more thoroughly address this fairness issue. For instance, one possible direction is to identify Pareto-optimal designs that reduce overall review noise while maintaining a reasonable degree of fairness.
4.1 The Necessity of Partition-based Isotonic Mechanisms
{tblr}
colspec = ccccc, cell22 = red!30, cell32 = red!30, cell43 = yellow!30, cell44 = yellow!30, Item Owner & $1$ $2$ $3$
$1$ 1 0 1
$2$ 1 1 0
$3$ 0 1 1
  {tblr}colspec = ccccc, cell22 = red!30, cell32 = red!30, cell33 = yellow!30, cell43 = yellow!30, cell24 = cyan!20, cell44 = cyan!20, Item Owner & $1$ $2$ $3$
$1$ 1 0 1
$2$ 1 1 0
$3$ 0 1 1
Table 3: An illustration of partition-based mechanism (left) and a more general mechanism (right). In the partition-based mechanism, each color denotes a partitioned item set block. In the more general mechanism, different owners can have different item set partitions, as marked by different colors.
At this point, one may wonder whether restricting ranking elicitation to partitioned item sets is necessary, as it inevitably forces the mechanism to ignore some of the owners’ revealed information across the partition blocks. For example, consider an instance with $m=n=3,\mathcal{I}^{1}=\{1,2\},\mathcal{I}^{2}=\{2,3\},\mathcal{I}^{3}=\{1,3\}$ — i.e., each owner owns two items, and each item has two co-owners (see the illustration in Table 3). We can see that, with any partition of the item set, Mechanism 2 can elicit ranking information from at most one owner. However, a more aggressive design could be eliciting the ranking information from every owner $j$ about her items in $\mathcal{I}^{j}$ and then applying a generalized version of Mechanism 1 by using the elicited (partial) ranking of $j$’s items. This is certainly a statistically more efficient design, but the key question is whether a mechanism of such kind can still be guaranteed to be truthful. Our study next shows that the answer is unfortunately “No”. We show that, within a much broader class of isotonic-regression-based mechanisms, the partition-based mechanism as prescribed in Mechanism 2 is essentially the only candidate that can guarantee truthful owner behaviors — that is, one may have to give up eliciting the comparison information between the partitioned blocks in order to trade for incentive properties. This reveals an intrinsic tradeoff between statistical efficiency and incentive guarantee within the general class of Isotonic Mechanisms. We leave it as an intriguing open direction to explore alternative estimation methods other than isotonic regression that can balance statistical efficiency and incentive compatibility — no such method is known so far even for the single-owner case.
We start by generalizing Mechanism 2 to a broader class of Isotonic Mechanisms, as described in Mechanism 3. In this more generalized Isotonic Mechanism, the calibrated score of any item $i\in[n]$ is now allowed to depend on the information elicited from all its owners, denoted by $\mathcal{J}^{i}\subseteq[m]$. Concretely, the input to Mechanism 2 and Mechanism 3 is the same, but the parameter choices in Mechanism 3 are strictly more general: (1) the item partition $\mathfrak{S}^{j}=\{{\mathcal{S}}^{j}_{1},{\mathcal{S}}^{j}_{2},\dots\mathcal{S}^{j}_{K^{j}}\}$, satisfying $\bigcup_{k=1}^{K^{j}}{\mathcal{S}}_{k}^{j}=\mathcal{I}^{j},{\mathcal{S}}_{k}^{j}\cap{\mathcal{S}}_{k^{\prime}}^{j}=\varnothing$, is now allowed to be different across owners; (2) additionally, a weight vector $\bm{\beta}^{j}=[\beta_{1}^{j},\beta_{2}^{j},\dots,\beta_{n}^{j}]\in\mathbb{R}_{\geq 0}^{n}$ is introduced to allow owner $j$’s more fine-grained influence on her items with itemized weights. This is much more powerful than $j$’s influence in Mechanism 2 that is restricted to be the same among items. The parameters are also determined from the input problem instance, specifically the ownership relation $\{\mathcal{I}^{j}\}_{j=1}^{m}$. The adjusted score for each item is then similarly determined by a weighted linear combination of the rank-calibrated scores from each owner. Notably, parameter $\beta_{i}^{j}$ for any $i\not\in\mathcal{I}^{j}$ is never used by Mechanism 3 hence can be arbitrary. We nevertheless defined $\bm{\beta}^{j}$ as a vector in $\mathbb{R}_{\geq 0}^{n}$ mainly for notational convenience.
Parameters : Personalized item partition $\mathfrak{S}^{j}=\{{\mathcal{S}}_{1}^{j},\ldots,{\mathcal{S}}_{K^{j}}^{j}\}$ and itemized weights $\bm{\beta}^{j}$ for each owner $j\in[m]$ based on ownership relations $\{\mathcal{I}^{j}\}_{j=1}^{m}$.
Input: Review scores $\bm{y}\in\mathbb{R}^{n}$, owner credentials $\{\alpha^{j}\}_{j=1}^{m}$.
for every $j\in[m]$ do
       Apply Mechanism 1 to owner $j$ with partition $\mathfrak{S}^{j}$ to estimate the ground-truth score, denoted by $\widehat{R}^{j}_{i}$, for each item $i$ in $\mathcal{I}^{j}$.
return $\big{\{}\widehat{R}_{i}=\sum_{j\in\mathcal{J}^{i}}\alpha^{j}\beta^{j}_{i}\widehat{R}^{j}_{i}/\sum_{{j^{\prime}}\in\mathcal{J}^{i}}\alpha^{j^{\prime}}\beta^{j^{\prime}}_{i}\big{\}}_{i=1}^{n}$.
Mechanism 3 Isotonic Mechanism with Personalized Partition
It is easy to see that the class of Mechanism 3 strictly contains the class of Mechanism 2. On the one hand, it can be immediately verified that Mechanism 2 with any item set partition $\mathfrak{S}=\{{\mathcal{S}}_{1},\ldots,{\mathcal{S}}_{K}\}$ can be formulated as Mechanism 3 with some parameter $\big{\{}\mathfrak{S}^{j},\bm{\beta}^{j}\big{\}}_{j=1}^{m}$ (see the proof of Theorem 2 for the details) On the other hand, Mechanism 3 does have strictly more expressivity than Mechanism 2. For example, the non-partition-based mechanism illustrated above in Table 3 can be captured by the following elicitation rule: $\mathfrak{S}^{1}=\{\{1,2\}\},\bm{\beta}^{1}=[1,1,0],\mathfrak{S}^{2}=\{\{2,3\}\},\bm{\beta}^{2}=[0,1,1]$ and so on. One may now wonder whether this strictly more general Mechanism 3 is also strictly more powerful. Our next result shows that the answer is “NO” if the mechanism is truthful.
Theorem 2 (The Necessity of Global Partition).
Under Assumptions 2.1, 2.2, and 2.3, for any $\mathcal{M}$ in the format of Mechanism 3 that is truthful for every input, there exists a $\mathcal{M}^{\prime}$ in the format of Mechanism 2 that elicits no less ranking information. Formally, if the order of any two items $i,i^{\prime}$ is truthfully elicited by $\mathcal{M}$, then their order is also truthfully elicited by $\mathcal{M}^{\prime}$.
The formal proof is somewhat involved hence deferred to Appendix D. At a high level, our proof is structured with two main steps. First, we characterize the necessary conditions under which Mechanism 3 is truthful, which is based on the following key Lemma 1. Second, we show that for any Mechanism 3 under the necessary condition, a Mechanism 2 can be constructed to elicit at least as much ranking information.
Lemma 1.
Under Assumptions 2.1, 2.2, and 2.3, if a Mechanism 3 with parameters $\big{\{}\mathfrak{S}^{j},\bm{\beta}^{j}\big{\}}_{j=1}^{m}$ is truthful, then the following two conditions must both hold:
  1. (I)
    Each owner has balanced influence on the items within each partition blocks for any input, i.e.,
    for $\text{for any }j\in[m],{\mathcal{S}}\in\mathfrak{S}^{j},i,i^{\prime}\in{\mathcal{S}},\qquad\text{we have }\,\omega^{j}_{i}=\omega^{j}_{i^{\prime}},$
    where $\omega_{i}^{j}=\alpha^{j}\beta^{j}_{i}/\sum_{j^{\prime}\in\mathcal{J}^{i}}\alpha^{j^{\prime}}\beta^{j^{\prime}}_{i}$ denotes owner $j$’s relative influence on the score of item $i$.
  2. (II)
    The parameters $\big{\{}\mathfrak{S}^{j},\bm{\beta}^{j}\big{\}}_{j=1}^{m}$ has a valid partition structure in the following sense,
    for $\text{for any }j,j^{\prime}\in[m],{\mathcal{S}}\in\mathfrak{S}^{j},{\mathcal{S}}^{\prime}\in\mathfrak{S}^{j^{\prime}},\text{ if }{\mathcal{S}}\cap{\mathcal{S}}^{\prime}\neq\varnothing,\text{ then }\beta^{j}_{i}=\beta^{j}_{i^{\prime}},\text{ for all }i,i^{\prime}\in{\mathcal{S}}\cup{\mathcal{S}}^{\prime}.$ (1)
Moreover, these two conditions are equivalent to each other.